### v1

@@ 1 Introduction
This paper develops a general and model-free framework for fixed-income pricing based solely on static arbitrage. We show that the absence of (strict) arbitrage across a universe of fixed-income instruments is equivalent to the existence of a strictly positive (nonnegative) discount curve that reproduces all observed market prices. This result constitutes a fundamental theorem of fixed-income pricing and provides a theoretical foundation for discount-curve construction in insurance asset–liability management. To the best of our knowledge, this paper provides the first systematic treatment of static arbitrage in fixed-income price systems formulated directly in terms of dated cash flows and observed market prices, with an explicit focus on liability replication.
@@ 2 A Fundamental Theorem of Fixed-Income Pricing
We consider a universe of $M$ fixed-income instruments, indexed by $i=1,\dots,M$, with prices $P_{i}$ and deterministic cash flows $C_{i1},\dots,C_{iN}$ occurring at a common set of dates $0<x_{1}<\cdots<x_{N}$. [^4] If instrument $i$ has no cash flow at $x_{j}$, we set $C_{ij}=0$. We collect the prices in the $M\times 1$ vector $P$ and the cash flows in the $M\times N$ cash-flow matrix $C$.
We assume a frictionless market in which these instruments can be traded freely. A portfolio is represented by a vector $q\in\mathbb{R}^{M}$, where $q_{i}$ denotes the position in instrument $i$. Such a portfolio generates the cash-flow vector $q^{\top}C$ and has price $q^{\top}P$.
@@
the sets of cash-flow vectors generated by portfolios with strictly negative and nonpositive prices, respectively. By construction, $\mathcal{C}$ is an open convex cone in $\mathbb{R}^{N}$, $\overline{\mathcal{C}}$ is a closed convex cone, and indeed $\overline{\mathcal{C}}=\operatorname{cl}(\mathcal{C})$.

### v2

@@ 1 Introduction
This paper develops a general and model-free framework for fixed-income pricing based solely on static arbitrage. We show that the absence of (strict) arbitrage across a universe of fixed-income instruments is equivalent to the existence of a strictly positive (nonnegative) discount curve that reproduces all observed market prices. This result constitutes a fundamental theorem of fixed-income pricing and provides a theoretical foundation for discount-curve construction in insurance asset–liability management, as studied, for example, in [Filipovic2022, CamenzindFilipovic2024]. To the best of our knowledge, this paper provides the first systematic treatment of static arbitrage in fixed-income price systems formulated directly in terms of dated cash flows and observed market prices, with an explicit focus on liability replication.
@@ 2 A Fundamental Theorem of Fixed-Income Pricing
We consider a universe of $M$ fixed-income instruments, indexed by $i=1,\dots,M$, with prices $P_{i}$ and deterministic cash flows $C_{i1},\dots,C_{iN}$ occurring at a common set of dates $0<x_{1}<\cdots<x_{N}$. [^4] If instrument $i$ has no cash flow at date $x_{j}$, we set $C_{ij}=0$. We collect the prices in the vector $P\in\mathbb{R}^{M}$ and the cash flows in the $M\times N$ cash-flow matrix $C$. We also denote by $\bm{x}\coloneq[x_{1},\dots,x_{N}]^{\top}$ the vector of cash-flow dates, and for any function $g\colon[0,\infty)\to\mathbb{R}$ we write $g(\bm{x})\coloneq[g(x_{1}),\dots,g(x_{N})]^{\top}$.
We assume a frictionless market in which these instruments can be traded freely. A portfolio is represented by a vector $q\in\mathbb{R}^{M}$, where $q_{i}$ denotes the position in instrument $i$. Such a portfolio has price $q^{\top}P$ and generates the cash flows $q^{\top}C$.
@@
the sets of cash-flow vectors generated by portfolios with strictly negative and nonpositive prices, respectively. By construction, $\mathcal{C}$ is an open convex cone in $\mathbb{R}^{N}$, $\overline{\mathcal{C}}$ is a closed convex cone, and indeed $\overline{\mathcal{C}}$ is the closure of $\mathcal{C}$}.